OpenStudy (anonymous):

Why do 2 resistors in parallel together contain less resistance than the same 2 resistors in series?

OpenStudy (anonymous):

well i kinda have to explain through words

OpenStudy (anonymous):

Re(equivalent resistor) for R1 and R2 in parallel =(R1*R2)/(R1+R2) Re for series =R1+R2 sorry for the first answer

OpenStudy (anonymous):

because we have as a propriety for R1 and R1 two resistors : in parallel =====> 1/Re = (1/R1)+ (1/R2) in series ======> Re= R1+R2. /*Re is equivalent of the two*/

OpenStudy (anonymous):

mathematically if you take two positive numbers a and b , we always have : a*b/(a+b) <a+b

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

with pleasure

OpenStudy (anonymous):

For the same reason traffic flows easier on a highway with two lanes instead of one. You have another whole path for the electrons, so they flow easier.